OpenStudy (anonymous):

How long will it take for an investment of $3000 to double in value if the interest rate is 7.5% per year, compounded continuously? (Round your answer to two decimal places.)

OpenStudy (anonymous):

Vn=Or(1+or-rate)n Values after n years equals the original value times (1 +or- the rate)to the n years power

OpenStudy (anonymous):

6000=3000(1+.075)n

OpenStudy (anonymous):

solve for n and that's your answer

OpenStudy (anonymous):

is it 2790.6977 ?

OpenStudy (anonymous):

cause that doesnt make sense

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Actually, this is the equation because it's compounded continuously... 6000 = 3000e^(.075t) divide both sides by the 3000 and then take the natural log to get rid of the e. ln(2000) = .075t t = ln(2000)/.075 ~ 101.3453661... 101.35 years (rounded)

OpenStudy (anonymous):

ah i see thanks Ryan!